# Masks and Boxes: Combining the Best of Both Worlds for Multi-Object Tracking

### v2

@@ 4.1 Implementation details
For object detections, we use the YOLOX [15] detector pretrained on the relevant dataset, following the practices of our baseline [38], unless stated otherwise. Dataset-specific variations are detailed in the next subsection.
Detections are divided into high and low confidence sets using a threshold of 0.6 for all sequences. For KITTI-tracking (see Sec. 4.2), we adjust the threshold to 0.5 due to different detection quality. In both our baseline and McByte, the new tracklet initialization threshold is set 0.1 higher than the detection confidence threshold. Other parameters match those in the baseline [38].
For mask creation, we use SAM [20], ensuring fair comparison with related works, with the vit_b model and weights from SAM’s authors. As a mask temporal propagator, we use Cutie [6] with the weights provided by the authors, Cutie base mega.
We set the mask confidence threshold (condition 2) to 0.6, $mm_{2}$ threshold (condition 3) to 0.05, and $mm_{1}$ threshold (condition 4) to 0.9. All experiments are conducted on an Nvidia A100 GPU with 40 GB of memory.
@@ 4.2 Datasets and evaluation metrics
We evaluate McByte on four datasets with diverse characteristics: DanceTrack [33], MOT17 [28], SoccerNet-tracking 2022 [9], and KITTI-tracking [16], using different detection sources per dataset as per community practices.
@@
MOT17 [28] involves tracking people in public spaces under varying conditions, such as lighting, pedestrian density, and camera stability. YOLOX models, pre-trained by the baseline [38], are used for testing and ablation studies.
SoccerNet-tracking 2022 [9] contains soccer match videos, where players move rapidly and look alike in the same team. Camera movement is is always present, and oracle detections (confidence 1.0) are provided by the dataset authors, meaning all detections are treated as high confidence.
KITTI-tracking [16] captures pedestrian and vehicle scenes from a moving car, with varied pedestrian density and frequent camera movement. Following community practices [4, 12], we use PermaTr [34] detections and set a lower confidence threshold of 0.5 due to the different quality of these detections.
We report three standard MOT metrics: HOTA [26], MOTA [2], and IDF1 [31], focusing on HOTA and IDF1 for evaluating tracking performance, as MOTA primarily measures detection accuracy. Higher values in these metrics indicate better performance. Limited number of submissions is allowed on the test set evaluation servers of MOT17 [28] and KITTI-tracking [16].
@@ 4.3 Ablation studies
We perform an ablation study to demonstrate the impact of incorporating the mask as an association cue along with the conditions discussed in Sec. 3.3. We evaluate the following variants:
  • a1: Uses only the mask signal for association if the mask is visible for the given tracklet, without ambiguity checks (Sec. 3.3). The value of $mm^{i,j}_{2}$ is directly assigned to $costs^{i,j}$ in Eq. 3. No association occurs if there is no mask.
  • a2: Similar to a1, but if the mask is unavailable, intersection over union (IoU) scores are used for association, as in the baseline [38].
  • a3: Adds an ambiguity check (Sec. 3.3). If the mask is visible, mask and bounding box information are fused as shown in Eq. 3. If no mask is available, IoU scores are used as in the baseline.
@@
Method HOTA MOTA IDF1
basline [38]: no mask 47.1 88.2 51.9
a1: either mask or no assoc. 48.6 80.8 44.4
a2: either mask or IoU for assoc. 45.3 82.2 41.5
a3: IoU and mask if ambiguity 56.6 89.5 57.0
a4: a3 + mask confidence 57.3 89.6 57.7
a5: a4 + $mm_{2}$ 58.8 89.6 60.1
a6: a5 + $mm_{1}$ 62.1 89.7 63.4
McByte: a6 + cmc 62.3 89.8 64.0
Table 1: Ablation study on DanceTrack [33] validation set listing the effects of the imposed constraints on using the mask as an association cue.
Method HOTA MOTA IDF1
Baseline, DanceTrack val 47.1 88.2 51.9
McByte, DanceTrack val 62.3 89.8 64.0
Baseline, SoccerNet-tracking 2022 test 72.1 94.5 75.3
McByte, SoccerNet-tracking 2022 test 85.0 96.8 79.9
Baseline, MOT17 val 68.4 78.2 80.2
McByte, MOT17 val 69.9 78.5 82.8
Baseline, KITTI-tracking test 54.3 63.7 -
McByte, KITTI-tracking test 55.1 66.9 -
Table 2: Ablation study comparing McByte with the baseline [38] on four different datasets.
Method HOTA MOTA IDF1
ByteTrack [38] 47.7 89.6 53.9
OC-SORT [4] 55.1 92.2 54.9
Deep OC-SORT [27] 61.3 92.3 61.5
C-BIoU [35] * 45.8 88.4 52.0
StrongSORT++ [12] 55.6 91.1 55.2
Hybrid-SORT [36] 65.7 91.8 67.4
McByte (ours) 67.1 92.9 68.1
@@
Method HOTA MOTA IDF1
@@
ByteTrack [38] 63.1 80.3 77.3
StrongSORT++ [12] 64.4 79.6 79.5
OC-SORT [4] 63.2 78.0 77.5
Deep OC-SORT [27] 64.9 79.4 80.6
Hybrid-SORT [36] 64.0 79.9 78.7
@@
C-BIoU [35] * 62.4 79.5 77.1
McByte (ours) 64.2 80.2 79.4
@@
Method HOTA MOTA IDF1
ByteTrack 72.1 94.5 75.3
OC-SORT [4] 82.0 98.3 76.3
C-BIoU [35] * 72.7 95.4 76.4
McByte (ours) 85.0 96.8 79.9
@@
Table 6: Comparing McByte with state-of-the-art tracking-by-detection algorithms on KITTI-tracking test set [16].
@@
The results of each variant are listed Tab. 1. In variant a1, where only the mask signal is used for association, we can see that despite HOTA increase, IDF1 decrease is present with respect to the baseline. It is caused by the fact that the mask use is uncontrolled and chaotic. With mask possibly providing incorrect results, the association cues can be misleading. If we perform the association either based only on mask or only on IoU (depending on the availability of the mask), as in variant a2, we might face an inconsistency of the cues between tracklets and detections from the same frame as well as over the next frames. This might lead to performance degradation. However, when we use properly both cues fusing the available information (variant a3), we can observe significant performance gain. We explain it that the algorithm is initially designed to work on bounding boxes while mask is a valuable guiding cue which can enforce existing association mechanisms. When the mask signal is the only cue or not properly fused with the IoU cue, the lower performance might be obtained (as in a1 and a2).
@@
Adding the conditional check based on mask confidence (variant a4) further improves the performance, because sometimes mask might be uncertain or incorrect providing misleading association guidance. Adding the minimal $mm_{2}$ value check (variant a5) also provides performance gain, because this check filters out the tracklet masks which could be considered as a noise or tiny parts of people almost entirely occluded. Another performance gain can be observed with the minimal $mm_{1}$ value check (variant a6). This check determines if the mask of the tracked person is actually within the bounding box and not too much outside it. Since the detection box might not be perfect, we allow small part of the tracklet mask to be outside the bounding box, but its major part must be within the bounding box, so that the mask can be used for guiding the association between the considered tracklet-detection pair. As it is shown, it further helps. Finally, we add the camera motion compensation (CMC), denoted as McByte in Tab. 1, which also provides some performance gain. Since the camera is moving in some sequences, CMC improves the quality of the estimated bounding boxes.
@@
We compare our McByte tracking algorithm to the baseline [38] across the four datasets, as shown in Tab. 2. Performance improvements are seen in all cases, though the gains vary by dataset due to differing characteristics. For example, DanceTrack [33] features non-linear motion, where the mask signal is particularly helpful in tracking occluded subjects, whereas IoU might struggle. In SoccerNet-tracking 2022 [9], where occlusions are fewer but motion is more abrupt (e.g., players running), the mask aids in successfully capturing players, despite similar outfits among team members.
@@
The performance gain on MOT17 [28] is smaller than on the previous datasets due to crowded scenes and smaller, distant people, making it harder for the mask to capture them accurately. The quality of detections, especially for small or uncertain bounding boxes, makes it difficult for the mask to comply with our conditions ($mm_{1}$ and $mm_{2}$), which rely on bounding boxes. KITTI-tracking shows a similar trend, with fewer pedestrians but larger, more easily detectable cars. As shown in Tab. 6, better detection quality for cars leads to higher performance gains. Although gains vary by dataset, McByte consistently improves performance (Tab. 2), demonstrating its robustness and general applicability.
@@
MyByte also performs better than the baseline when the public detections of a different quality are given. Due to the space limits, we show it in the supplementary material, Appendix E.
@@ 4.4 Comparison with state of the art tracking-by-detection methods
We compare McByte with state-of-the-art tracking-by-detection algorithms across the test sets of the four datasets. The results, listed in Tabs. 3, 4, 5 and 6, show that McByte achieves the highest HOTA and IDF1 scores on DanceTrack [33] (Tab. 3) and SoccerNet-tracking 2022 [9] (Tab. 5), as well as the highest HOTA scores on KITTI-tracking [16] (Tab. 6).
The KITTI-tracking test server does not provide IDF1 scores and evaluates both pedestrian and car classes. We also test ByteTrack [38] with the same detection set for comparison. ByteTrack performs poorly on the car class because it is designed for tracking people and is optimized for its pre-trained detections. When using PermaTr [34] detections, which include cars, ByteTrack struggles to generalize, leading to lower performance, though it performs comparably on the pedestrian class.
On the MOT17 [28] test set, unlike the baseline [38] and derived methods [12, 4, 27, 36], we do not tune parameters per sequence. With our mask-based association cue, we achieve improvements over the baseline and comparable performance with other methods, despite not tuning parameters. Due to limited evaluations allowed on the MOT17 test server [10], we are unable to evaluate all listed tracking-by-detection methods without sequence-specific tuning.
Additionally, C-BIoU [35] is marked with an asterisk in Tabs. 3, 4 and 5. As no implementation is published, and not all necessary details for reproduction are provided, we reproduce the method based on the available descriptions and report the best results we have obtained. We descibe it more in detail in the supplementary material, Appendix C.
@@ 4.5 Comparison with other methods using mask
Grounded SAM 2 [29, 24] struggles with segment-based tracking, where merging segments into full tracklets is unreliable. Though YOLOX detections improve the performance, the results remain unsatisfactory and far below McByte’s performance.
@@
These results highlight that existing mask-based methods are unsuitable for MOT. In contrast, McByte effectively combines mask association with bounding box processing and tracklet management, making it better suited for MOT tasks.
More results and details on DEVA, Grounded SAM 2, and MASA experiments are available in the supplementary material, Appendix A
Method HOTA MOTA IDF1
DEVA, original settings 31.8 -89.4 31.3
DEVA, with YOLOX 24.7 -239.7 20.4
Grounded SAM 2, original settings 43.4 18.4 47.6
Grounded SAM 2, with YOLOX 47.5 43.0 54.1
MASA, original settings 45.5 36.9 53.6
MASA, with YOLOX 63.5 74.0 73.6
McByte (ours) 69.9 78.5 82.8
@@ 5 Conclusion
In this paper, we explore the use of a temporally propagated segmentation mask as an association cue for MOT. We develop a new mechanism that incorporates the mask into a tracking-by-detection approach, combining mask and bounding box information to improve tracking performance. Results from four datasets show the effectiveness of this approach, with our algorithm boosting MOT performance. Additionally, it outperforms other systems that rely primarily on segmentation masks, highlighting that when carefully managed, the mask can serve as a powerful association cue for MOT.
@@ Appendix A More experiments and details with mask-based tracking systems
Details HOTA MOTA IDF1
@@
GDino ”person”, th. 0.35 $\ddagger$ 31.8 -89.4 31.3
YOLOX ByteTrack, th. 0.6 $\ddagger$ 24.7 -239.7 20.4
YOLOX ByteTrack, th. 0.7 27.0 -187.8 23.7
@@
GDino ”person”, th. 0.25, step 20 $\ddagger$ 43.4 18.4 47.6
GDino ”person”, th. 0.25, step 100 44.0 15.5 49.0
YOLOX ByteTrack, th. 0.25, step 20 46.4 36.0 51.6
YOLOX ByteTrack, th. 0.6, step 20 $\ddagger$ 47.5 43.0 54.1
YOLOX ByteTrack, th. 0.7, step 20 47.4 44.3 54.1
YOLOX ByteTrack, th. 0.25, step 100 46.8 30.2 54.2
YOLOX ByteTrack, th. 0.6, step 100 47.4 34.8 54.9
YOLOX ByteTrack, th. 0.7, step 100 47.4 35.9 54.9
YOLOX ByteTrack, th. 0.25, step 1 43.0 36.2 43.9
YOLOX ByteTrack, th. 0.6, step 1 44.4 44.9 46.5
YOLOX ByteTrack, th. 0.7, step 1 44.3 46.5 46.7
@@
GDino feat. Detic-SwinB ”person”, th 0.2 46.8 24.3 52.1
GDino feat. YOLOX COCO, th 0.3 45.4 36.9 53.1
GDino feat. YOLOX ByteTrack, th 0.3 61.8 71.3 70.8
GDino feat. YOLOX ByteTrack, th 0.6 63.4 73.8 73.3
GDino feat. YOLOX ByteTrack, th 0.7 62.5 72.9 71.9
R50 feat. YOLOX COCO, th 0.3 $\ddagger$ 45.5 36.9 53.6
R50 feat. YOLOX ByteTrack, th 0.3 62.5 71.5 72.0
R50 feat. YOLOX ByteTrack, th 0.6 $\ddagger$ 63.5 74.0 73.6
R50 feat. YOLOX ByteTrack, th 0.7 62.6 73.0 72.3
@@
Method HOTA MOTA IDF1
DEVA, original settings 21.9 -347.1 15.8
DEVA, with YOLOX 20.1 -423.9 13.3
Grounded SAM 2, original settings 51.3 73.5 48.0
Grounded SAM 2, with YOLOX 52.9 81.6 49.6
MASA, original settings 38.2 71.9 34.9
MASA, with YOLOX 46.0 85.6 41.1
McByte (ours) 62.3 89.8 64.0
@@
We evaluate DEVA [7], Grounded SAM 2 [29, 24], and MASA [22] on MOT datasets, saving each bounding box output per frame in MOT format [10]. Grounded SAM 2 and MASA provide object IDs, while for DEVA, we use unique, immutable IDs from the propagated masks associated with each bounding box.
@@
For DEVA, we first run the default settings using the Grounding Dino [24] detector with the ”person” prompt and a confidence threshold of 0.35 to accept bounding boxes. Then, we replace it with the YOLOX [15] detector, trained on the MOT17 dataset from our baseline [38]. We test two threshold values, 0.6 and 0.7, which correspond to the high-confidence detection threshold and the new tracklet initialization in our method, respectively (Sec. 4.1 in the main paper).
@@
Next, we integrate YOLOX detector with weights from our baseline [38] and run variants with step values of 20, 100, and 1, using different bounding box allowance thresholds of 0.25, 0.6, and 0.7 (analogous to the DEVA experiments). We also attempt to run a variant with the segment length set to the entire video sequence, but it fails due to excessive GPU memory requirements. Additionally, this setup would only track objects visible in the first frame.
@@
We also incorporate the YOLOX detector with weights from our baseline [38] and test variants with detection confidence thresholds of 0.3, 0.6, and 0.7, analogously to DEVA and Grounded SAM 2. Additionally, we run the ResNet-50 feature variants with the YOLOX COCO model (threshold 0.3) and the baseline-pre-trained weights (thresholds 0.3, 0.6, 0.7).
@@
Tab. 9 presents the performance of DEVA, Grounded SAM 2, and MASA on the DanceTrack [33] validation set. The listed variants correspond to those marked with $\ddagger$ in  Tab. 8 and are the ones reported in the main paper.
@@
Appendix B State-of-the-art comparison with offline and transformer-based methods
Method HOTA MOTA IDF1

### v3

@@ 4.1 Implementation details
For object detections, we follow the practices of our baseline [40] and use the YOLOX [15] detector pretrained on the relevant dataset, unless stated otherwise. As in baseline, detections are divided into high and low confidence sets based on a defined threshold. However, on the contrary to the baseline which considers different values per video sequence, we always use the same threshold for all sequences. We fix it as 0.6, which is the default value of our baseline, i.e. when the baseline is evaluated on datasets without tuning per sequence.
For mask creation, we use SAM [20], ensuring fair comparison with related works, with the vit_b model and weights from SAM’s authors. As a mask temporal propagator, we use Cutie [8] with the weights provided by the authors, Cutie base mega. Note that we use the image segmentation model (SAM) only to create and initialize the masks for the newly appeared objects at the scene. For the association cue between the tracklets and detections we process the propagated mask from the mask temporal propagator (Cutie). All experiments are conducted on an Nvidia A100 GPU with 40 GB of memory.
@@ 4.2 Datasets and evaluation metrics
We evaluate McByte on four person tracking datasets with diverse characteristics, to demonstrate the generalizability of our method. We present results on DanceTrack [33], MOT17 [28], SoccerNet-tracking 2022 [9], and KITTI-tracking [16], while using detection sources per dataset as per community practices for fair comparison.
@@
MOT17 [28] involves tracking people in public spaces under varying conditions, such as lighting, pedestrian density, and camera stability. We use YOLOX pre-trained on MOT17 for detections as provided by the the baseline [40].
SoccerNet-tracking 2022 [9] contains soccer match videos, where players move rapidly and look alike in the same team. Camera movement is always present and oracle detections are provided by the dataset authors.
KITTI-tracking [16] captures scenes from car view, with varied pedestrian density and frequent camera movement. It also considers another class, car, which can help to assess the generalizability of the tracking method. Following community practices [4, 12], we use PermaTr [34] detections.
We report three standard MOT metrics: HOTA [26], IDF1 [31] and MOTA [2], focusing on HOTA and IDF1 for evaluating the tracking performance. IDF1 measures the tracking quality and identity preservation, while HOTA includes association, detection and localization. MOTA primarily measures detection quality and we report it for the result completeness. Higher values in these metrics indicate better performance.
Method HOTA IDF1 MOTA
basline [40]: no mask 47.1 51.9 88.2
a1: either mask or no assoc. 48.6 44.4 80.8
a2: either mask or IoU for assoc. 45.3 41.5 82.2
a3: IoU and mask if ambiguity 56.6 57.0 89.5
a4: a3 + mask confidence 57.3 57.7 89.6
a5: a4 + $mm_{2}$ 58.8 60.1 89.6
a6: a5 + $mm_{1}$ 62.1 63.4 89.7
McByte: a6 + cmc 62.3 64.0 89.8
Table 1: Ablation study on DanceTrack [33] validation set listing the effects of the imposed constraints on using the temporally propagated mask as an association cue.
Method HOTA IDF1 MOTA
Baseline, DanceTrack val 47.1 51.9 88.2
McByte, DanceTrack val 62.3 64.0 89.8
Baseline, SoccerNet-tracking 2022 test 72.1 75.3 94.5
McByte, SoccerNet-tracking 2022 test 85.0 79.9 96.8
Baseline, MOT17 val 68.4 80.2 78.2
McByte, MOT17 val 69.9 82.8 78.5
Baseline, KITTI-tracking test 54.3 - 63.7
McByte, KITTI-tracking test 57.0 - 68.9
Table 2: Ablation study comparing McByte with the baseline [40] on four different datasets. As SoccerNet-tracking 2022 [9] and KITTI-tracking [16] (pedestrian) do not contain validation set split, we report the results on the test sets. KITTI evaluation server does not provide IDF1 scores.
@@ 4.3 Ablation studies
We perform an ablation study to demonstrate the impact of incorporating the temporally propagated (TP) mask as an association cue along with the conditions discussed in Sec. 3.3. We evaluate the following variants:
  • a1: Uses only the TP mask signal for association if the mask is visible for the given tracklet, without ambiguity checks (Sec. 3.3). The value of $mm^{i,j}_{2}$ is directly assigned to $costs^{i,j}$ in Eq. 3. No association occurs if there is no mask.
  • a2: Similar to a1, but if the TP mask is unavailable, intersection over union (IoU) scores are used for association, as in the baseline [40].
  • a3: Adds an ambiguity check (Sec. 3.3). If the TP mask is visible, mask and bounding box information are fused as shown in Eq. 3. If no mask is available, IoU scores are used as in the baseline.
@@
Method HOTA IDF1 MOTA
ByteTrack [40] 47.7 53.9 89.6
OC-SORT [4] 55.1 54.9 92.2
Deep OC-SORT [27] 61.3 61.5 92.3
C-BIoU [35] * 45.8 52.0 88.4
StrongSORT++ [12] 55.6 55.2 91.1
Hybrid-SORT [36] 65.7 67.4 91.8
McByte (ours) 67.1 68.1 92.9
@@
Method HOTA IDF1 MOTA
@@
ByteTrack [40] 63.1 77.3 80.3
StrongSORT++ [12] 64.4 79.5 79.6
OC-SORT [4] 63.2 77.5 78.0
Deep OC-SORT [27] 64.9 80.6 79.4
Hybrid-SORT [36] 64.0 78.7 79.9
@@
ByteTrack [5] 62.8 77.1 78.9
C-BIoU [35] * 62.4 77.1 79.5
McByte (ours) 64.2 79.4 80.2
@@
Method HOTA IDF1 MOTA
ByteTrack [40] 72.1 75.3 94.5
OC-SORT [4] 82.0 76.3 98.3
C-BIoU [35] * 72.7 76.4 95.4
McByte (ours) 85.0 79.9 96.8
@@
Table 6: Comparing McByte with state-of-the-art tracking-by-detection algorithms on KITTI-tracking test set [16]. KITTI evaluation server does not provide IDF1 scores.
@@
The results of each variant are listed in Tab. 1. In variant a1, where only the TP mask signal is used for association, we can see that despite HOTA increase, IDF1 decreases with respect to the baseline. It is caused by the fact that the mask use is uncontrolled and chaotic. With TP mask possibly providing incorrect results, the association cues can be misleading. If we perform the association either based only on TP mask or only on IoU (depending on the availability of the mask), as in variant a2, we might face an inconsistency of the cues between tracklets and detections from the same frame and the next frames. This might lead to performance degradation. However, when we use properly both cues fusing the available information (variant a3), we can observe significant performance gain. We explain it as the algorithm is initially designed to work on bounding boxes while TP mask is a valuable guiding cue which can improve existing association mechanisms. When the mask signal is the only cue or not properly fused with the IoU cue, a lower performance might be obtained (as in a1 and a2).
@@
Adding the conditional check based on TP mask confidence (variant a4) further improves the performance, because sometimes mask might be uncertain or incorrect providing misleading association guidance. Adding the minimal $mm_{2}$ value check (variant a5) also provides performance gain, because this check filters out the tracklet TP masks which could be considered as a noise or tiny parts of people almost entirely occluded. Another performance gain can be observed with the minimal $mm_{1}$ value check (variant a6). This check determines if the TP mask of the tracked person is actually within the bounding box and not too much outside it. Since the detection box might not be perfect, we allow small parts of the tracklet mask to be outside the bounding box, but its major part must be within the bounding box, so that the TP mask can be used for guiding the association between the considered tracklet-detection pair. As it is shown, it further helps. Finally, we add the camera motion compensation (CMC), denoted as McByte in Tab. 1, which also provides some performance gain.
@@
We compare our McByte tracking algorithm to the baseline [40] across the four datasets, as shown in Tab. 2. Performance improvements are seen in all cases, though the gains vary by dataset due to different characteristics. For example, DanceTrack [33] features non-linear motion, where the TP mask signal is particularly helpful in tracking occluded subjects, whereas IoU might struggle. In SoccerNet-tracking 2022 [9], where occlusions are fewer but motion is more abrupt (e.g., players running), the mask aids in successfully capturing players, despite similar outfits among team members.
@@
The performance gain on MOT17 [28] is smaller than on the previous datasets due to crowded scenes and smaller, distant people, making it harder for the TP mask to capture them accurately. The quality of detections, especially for small or uncertain bounding boxes, makes it difficult for the mask to comply with our conditions ($mm_{1}$ and $mm_{2}$), which rely on bounding boxes. KITTI-tracking shows a similar trend yet with higher gain, with less distant pedestrians and different quality of detections. Although gains vary by dataset, McByte consistently improves performance (Tab. 2), demonstrating its robustness and general applicability. TP mask incorporated as an association cue helps to handle challenging scene situations such as high occlusions and reduced subject visibility as shown in Fig. 1.
@@
MyByte also performs better than the baseline when the public detections are given. Due to the space limits, we show it in Appendix D.
@@ 4.4 Comparison with state of the art tracking-by-detection methods
We compare McByte with state-of-the-art tracking-by-detection algorithms across the test sets of the four diversed datasets. The results, listed in Tabs. 3, 4, 5 and 6 show that McByte achieves the highest HOTA, IDF1 and MOTA scores on DanceTrack [33] (Tab. 3), MOT17 [28] (Tab. 4) and KITTI-tracking [16] (Tab. 6). In case of SoccerNet-tracking 2022 [9] (Tab. 5), McByte achieves the highest HOTA and IDF1 scores, and the second highest MOTA score.
The KITTI-tracking test server evaluates both pedestrian and car classes. We also test ByteTrack [40], our baseline, with the same detection set for comparison. ByteTrack performs poorly on the car class because its parameters are specifically set for tracking people. When using PermaTr [34] detections, which include cars, ByteTrack struggles to generalize, leading to lower performance, though it performs comparably on the pedestrian class. Our method performs well on pedestrians, but also on other objects, such as cars, which we demonstrate in Tab. 6.
On the MOT17 [28] test set, unlike the baseline [40] and derived methods [12, 4, 27, 36], we do not tune parameters per sequence. We aim at developing a generalizable tracker without the need for changes on different sequences or datasets. We reach the best scores among the non-tuned trackers. We use the result of ByteTrack [40] not being tuned per sequence as reported in [5]. Juxtaposed with the methods which do tune their parameters per sequence, we achieve improvements over the baseline [40] and comparable performance with other methods.
Additionally, C-BIoU [35] is marked with an asterisk in Tabs. 3, 4 and 5. As no implementation is published, and not all necessary details for reproduction are provided, we reproduce the method based on the available descriptions and report the best results we have obtained. We describe it more in detail in Appendix C.
@@ 4.5 Comparison with other methods using mask
Grounded SAM 2 [29, 24] struggles with segment-based tracking, where merging segments into full tracklets is unreliable. Though YOLOX detections improve the performance, the results remain far below McByte’s performance.
@@
These results highlight that existing mask-based methods are unsuitable for MOT. In contrast, McByte effectively combines temporally propagated mask-based association with bounding box processing and tracklet management, making it better suited for MOT tasks.
More results and details on DEVA, Grounded SAM 2, and MASA experiments are available in Appendix A.
Method HOTA IDF1 MOTA
DEVA, original settings 31.8 31.3 -89.4
DEVA, with YOLOX 24.7 20.4 -239.7
Grounded SAM 2, original settings 43.4 47.6 18.4
Grounded SAM 2, with YOLOX 47.5 54.1 43.0
MASA, original settings 45.5 53.6 36.9
MASA, with YOLOX 63.5 73.6 74.0
McByte (ours) 69.9 82.8 78.5
@@ 5 Conclusion
In this paper, we explore the use of a temporally propagated segmentation mask as an association cue for MOT. We develop a new mechanism that incorporates the mask propagation into a tracking-by-detection approach, fusing temporally propagated mask and bounding box information to improve bounding box-based multi-object tracking performance. Results from four datasets show the effectiveness of this approach, with our algorithm boosting MOT performance. Additionally, it outperforms other systems that rely primarily on segmentation masks, highlighting that when carefully managed, the mask can serve as a powerful association cue for MOT.
@@ Appendix A More experiments and details with mask-based tracking systems
Details HOTA IDF1 MOTA
@@
GDino ”person”, th. 0.35 $\ddagger$ 31.8 31.3 -89.4
YOLOX ByteTrack, th. 0.6 $\ddagger$ 24.7 20.4 -239.7
YOLOX ByteTrack, th. 0.7 27.0 23.7 -187.8
@@
GDino ”person”, th. 0.25, step 20 $\ddagger$ 43.4 47.6 18.4
GDino ”person”, th. 0.25, step 100 44.0 49.0 15.5
YOLOX ByteTrack, th. 0.25, step 20 46.4 51.6 36.0
YOLOX ByteTrack, th. 0.6, step 20 $\ddagger$ 47.5 54.1 43.0
YOLOX ByteTrack, th. 0.7, step 20 47.4 54.1 44.3
YOLOX ByteTrack, th. 0.25, step 100 46.8 54.2 30.2
YOLOX ByteTrack, th. 0.6, step 100 47.4 54.9 34.8
YOLOX ByteTrack, th. 0.7, step 100 47.4 54.9 35.9
YOLOX ByteTrack, th. 0.25, step 1 43.0 43.9 36.2
YOLOX ByteTrack, th. 0.6, step 1 44.4 46.5 44.9
YOLOX ByteTrack, th. 0.7, step 1 44.3 46.7 46.5
@@
GDino feat. Detic-SwinB ”person”, th 0.2 46.8 52.1 24.3
GDino feat. YOLOX COCO, th 0.3 45.4 53.1 36.9
GDino feat. YOLOX ByteTrack, th 0.3 61.8 70.8 71.3
GDino feat. YOLOX ByteTrack, th 0.6 63.4 73.3 73.8
GDino feat. YOLOX ByteTrack, th 0.7 62.5 71.9 72.9
R50 feat. YOLOX COCO, th 0.3 $\ddagger$ 45.5 53.6 36.9
R50 feat. YOLOX ByteTrack, th 0.3 62.5 72.0 71.5
R50 feat. YOLOX ByteTrack, th 0.6 $\ddagger$ 63.5 73.6 74.0
R50 feat. YOLOX ByteTrack, th 0.7 62.6 72.3 73.0
@@
Method HOTA IDF1 MOTA
DEVA, original settings 21.9 15.8 -347.1
DEVA, with YOLOX 20.1 13.3 -423.9
Grounded SAM 2, original settings 51.3 48.0 73.5
Grounded SAM 2, with YOLOX 52.9 49.6 81.6
MASA, original settings 38.2 34.9 71.9
MASA, with YOLOX 46.0 41.1 85.6
McByte (ours) 62.3 64.0 89.8
@@
We evaluate DEVA [7], Grounded SAM 2 [29, 24], and MASA [22] on MOT datasets, saving each bounding box output per frame in MOT format [10]. The limitations of these methods when applied to MOT are described in Secs. 2.3 and 4.5 of the main paper.
@@
For DEVA, we first run the default settings using the Grounding Dino [24] detector with the ”person” prompt and a confidence threshold of 0.35 to accept bounding boxes. Then, we replace it with the YOLOX [15] detector, trained on the MOT17 dataset from our baseline [40]. We test two threshold values, 0.6 and 0.7. In our baseline, initialization of the new tracklets happens for the values 0.1 higher than the high confidence detection threshold. As we consider the default value of 0.6 for the latter (Sec. 4.1 in the main paper), we also experiment with the value of 0.7 with DEVA and other mask based systems.
@@
Next, we integrate YOLOX detector with weights from our baseline [40] and run variants with step values of 20, 100, and 1 (thus processing detections every frame), using different bounding box allowance thresholds of 0.25, 0.6, and 0.7 (analogous to the DEVA experiments). We also attempt to run a variant with the segment length set to the entire video sequence, but it fails due to excessive GPU memory requirements. Additionally, this setup would only track objects visible in the first frame.
@@
Further, we incorporate the YOLOX detector with weights from our baseline [40] and test variants with detection confidence thresholds of 0.3, 0.6, and 0.7, analogously to DEVA and Grounded SAM 2. Additionally, we run the ResNet-50 feature variants with the YOLOX COCO model (threshold 0.3) and the baseline-pre-trained weights (thresholds 0.3, 0.6, 0.7).
@@
Tab. 9 presents the performance of DEVA, Grounded SAM 2, and MASA on the DanceTrack [33] validation set. The listed variants correspond to those marked with $\ddagger$ in  Tab. 8 and are the ones reported in the main paper on MOT17.
@@
Appendix B State-of-the-art comparison with transformer-based and other types of method
Method HOTA IDF1 MOTA